OpenStudy (anonymous):

how to do- \[\lim_{x \rightarrow 0} (\cos 2x -1)/(\cos x- 1 )\]

OpenStudy (kira_yamato):

Use l Hospital theorem (I don't know how to spell it) But it goes \[\lim_{x \rightarrow a} \left[ \frac{f(x)}{g(x)} \right] = \lim_{x \rightarrow a} \left[ \frac{f'(x)}{g'(x)} \right]\]

OpenStudy (zarkon):

\[\lim_{x \to0} \frac{\cos (2x) -1}{\cos (x)- 1}\] \[=\lim_{x \to0} \frac{\cos^2(x)-\sin^2(x) -1}{\cos (x)- 1}\] \[=\lim_{x \to0} \frac{-(1-\cos^2(x))-\sin^2(x) }{\cos (x)- 1}\] \[=\lim_{x \to0} \frac{-\sin^2(x)-\sin^2(x) }{\cos (x)- 1}\] \[=\lim_{x \to0} \frac{-2\sin^2(x)}{\cos (x)- 1}\] \[=\lim_{x \to0} \frac{-2\sin^2(x)}{\cos (x)- 1}\frac{\cos (x)+ 1}{\cos (x)+ 1}\] \[=\lim_{x \to0} \frac{-2\sin^2(x)}{\cos^2 (x)- 1}(\cos (x)+ 1)\] \[=\lim_{x \to0} \frac{-2\sin^2(x)}{-\sin^2 (x)}(\cos (x)+ 1)\] \[=\lim_{x \to0}\,\,\,\, 2(\cos (x)+ 1)=2(1+1)=4\]